The following exercise refers to choosing two cards from a thoroughly shuffled deck. Assume that the deck is shuffled after a card is returned to the deck. If you put the first card back in the deck before you draw the next, what is the probability that the first card is a 7 and the second card is a ace?
Answer #1

total cards = 52

probability that the first card is a 7 and the second card is a ace = 4/52 * 4/52

= 1/169 = 0.0059